# Multi-scale Restoration of Missing Data in Optical Time-series Images with Masked Spatial-Temporal Attention Network

### v1

@@ III-B3 Positional Encoding
In the Transformer, positional encoding is added to the input sequence to introduce spatio-temporal positional semantics, as shown in Equation (12), where $\operatorname{PosEnc}(pos,dim)$ represents the value of the $dim$-th dimension of the positional encoding for the $pos$-th element in the sequence.
@@
After positional encoding, the input sequence will undergo $L$ layers of MSTA to explore deeper spatio-temporal features. Finally, the overall expression of the MFE module is shown in Equation (13), where the symbol L denotes stacking $L$ layers, and $\mathrm{MFE}()$ represents the MSTA-based Feature Extractor.
@@ III-C Multi-scale Restoration Network
The MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) is a multi-scale restoration network composed of $S$ residual-connected restoration modules and observed value replacement, as illustrated in Fig. 4. Each module consists of Patch Embedding, MFE, and Patch Unembedding. These restoration modules can capture the spatio-temporal correlations of pixels at different scales to predict missing values and ultimately replace observed values as outputs.
@@ III-C1 Embedding and Unembedding
The input to MS2TAN consists of two components: the image sequence $X_{\mathrm{in}}$ and the missing value hint tensor $M$. To incorporate the distributional properties of the missing values into the model learning process, $X_{\mathrm{in}}$ and $M$ are concatenated along the channel dimension to yield the merged input $X_{c}=\mathrm{Concat}(X_{\mathrm{in}},M)$, where $X_{c}=\left[x_{c}^{1},x_{c}^{2},\cdots,x_{c}^{T}\right]\in\mathbb{R}^{T\times C^{\prime}\times H\times W}$, where $C^{\prime}=C+1$ is the total number of channels after concatenation.
Following the approach of Vision Transformer (ViT), each image is partitioned into a series of patches, with the size of each patch determined by the hyperparameter $P$. For each temporal image, $N=HW/P^{2}$ patches are obtained. These patches are then aggregated and reorganized to form a patch sequence of length $TN$, with each patch represented as a $C^{\prime}P^{2}$-dimensional vector. Equation (14) delineates the processing of the $t$-th temporal image $x_{c}^{t}$.
$\begin{gathered}A^{t}=\left[p_{1}^{t},p_{2}^{t},\cdots,p_{N}^{t}\right]=\operatorname{Patchify}(x_{c}^{t},P)\\ \text{where}\ x_{c}^{t}\in\mathbb{R}^{C^{\prime}\times H\times W}\ \text{and}\ A^{t}\in\mathbb{R}^{N\times C^{\prime}P^{2}}\end{gathered}$ (14)
Next, a linear layer is utilized to project the vectors corresponding to each patch into a high-dimensional representation space. The images from different time steps are then aggregated to obtain the complete token sequence $E$, as depicted in Equation (15).
@@
The process of transforming $X_{\mathrm{in}}$ and $M$ step by step to obtain $E$ as described above is referred to as Embedding, which can be represented by Equation (16). Conversely, if this operation is reversed, i.e., $E$ is restored to a patch sequence and then rearranged into an image, it is referred to as Unembedding, as shown in Equation (17). Fig. 5 illustrates the processes of Embedding and Unembedding on images.
$E=\mathrm{Embedding}\left(\mathrm{Concat}\left(X_{\mathrm{in}},M\right),P\right)$ (16)
@@ III-C2 MS2TAN
MS2TAN consists of components such as Embedding, MFE, and Unembedding, where the block size $P$ is an important hyperparameter, and the embedding scale of the $i$-th layer is denoted as $P^{\left(i\right)}$. The process of obtaining the embedding vector $\alpha$ from the input is described in Equation (18). Subsequently, MFE is utilized to mine deep spatio-temporal correlation features $\beta$, as shown in Equation (19). Finally, the features are decoded, unembedded back into image form, and added to the input, resulting in an intermediate result as depicted in Equation (20), where $\widetilde{Y}^{(i)}$ represents the $i$-th intermediate result.
$\alpha=\mathrm{Embedding}\left(\mathrm{Concat}\left(\widetilde{Y}^{(i-1)},M\right),P^{\left(i\right)}\right)$ (18)
@@
Let $\widetilde{Y}^{(0)}=X_{\mathrm{in}}$, and sequentially obtain $S$ intermediate results of the reconstruction. Eventually, the $S$-th result is taken as the output of MS2TAN, as shown in Equation (21).
@@ III-C3 Observed Value Replacement
For the original reconstruction results of this network, denoted as $\widetilde{Y}$, the observed values are replaced with their non-missing parts to obtain the final result $\widetilde{Y}_{\mathrm{out}}$, as shown in Equation (22).
$\widetilde{Y}_{\mathrm{out}}=\widetilde{Y}\odot(1-M)+X\odot M$ (22)
@@ III-D1 Pixel-wise Loss
Pixel-wise loss disregards the overall integrity of the image, treating the image as a collection of pixels, and comparing pixel by pixel to generate the image against the target image. It serves as the foundational loss for remote sensing image reconstruction tasks. Specifically, Equation (23) illustrates its calculation process.
@@
The pixel-wise loss calculated here includes both missing and observed parts. The reduction of the loss in the observed part does not directly improve the performance of the model. However, it does smooth the model output, making the model training process more stable, and hence is also included.
@@ III-D2 Structural Loss
Pixel-wise reconstruction loss is commonly employed in various time-series and visual tasks but fails to consider the correlation between pixels. Structural reconstruction loss, based on the principle of structural similarity [45], measures the difference with the target image using the structural similarity index (SSIM), optimizing the visual consistency of the reconstruction results from the perspectives of structure, contrast, and luminance.
@@
Equation (24) specifies the computation process of structural loss, where $\mu_{\eta}$ and $\mu_{y}$ denote the means of $\eta$ and $y$ respectively, $\sigma_{\eta}$ and $\sigma_{y}$ denote the variances of $\eta$ and $y$ respectively, $\sigma_{\eta y}$ denotes the covariance between $\eta$ and $y$, and $C_{1}$ and $C_{2}$ are constants.
@@ III-D3 Perceptual Loss
Perceptual loss utilizes a pre-trained VGG16 network [46], as the feature extraction network to obtain perceptual feature vectors, followed by applying L2 loss to the feature vectors. The calculation of perceptual loss is shown in Equation (25), where $\psi\left(\right)$ represents the feature extraction network, and $d_{f}$ denotes the feature dimension.
${\mathcal{L}}_{\mathrm{perceptual}}=\frac{1}{d_{f}}\left|\left|\psi\left(\eta\right)-\psi\left(y\right)\right|\right|_{2}^{2}$ (25)
@@ III-D4 Multi-Objective Loss Function
For each intermediate output $\widetilde{Y}^{(i)}$ of MS2TAN, the expression of the multi-objective loss function is shown in Equation (26), where $\lambda_{\mathrm{s}}$ and $\lambda_{\mathrm{p}}$ denote the weights of the structural loss and perceptual loss respectively. The final loss function considers the intermediate results of each layer, as shown in Equation (27).
@@
$\mathcal{L}^{\left(i\right)}\left(\eta,y\right)=\mathcal{L}_{\text{pixel-wise}}+{\lambda_{\mathrm{s}}\mathcal{L}}_{\mathrm{struct}}+{\lambda_{\mathrm{p}}\mathcal{L}}_{\mathrm{perceptual}}$ (26)
$\mathcal{L}=\frac{1}{S}\sum_{i=1}^{m}\mathcal{L}^{\left(i\right)}\left(\widetilde{Y}^{(i)},Y\right)$ (27)
Figure 6: The locations of the research areas used to evaluate the proposed methodology are shown with examples. S1 to S6 correspond to the locations of the six research areas in Table I: Wuhan, Hengyang, Shenzhen, Dallas, Kern, and Urumqi.
TABLE I: Details of Six Research Areas
  Satellite Task Location Date Range Area Size and Resolution
  S1 Landsat-5 TM ETM+ SLC-off Wuhan, Hubei, China 2001-03 to 2003-10 (1720, 2040) , 30-m
S2 Landsat-5 TM ETM+ SLC-off Hengyang, Hunan, China 2003-04 to 2005-05 (1200, 1200) , 30-m
S3 Landsat-5 TM Thick Cloud Removal Shenzhen, Guangdong, China 2003-10 to 2004-11 (1680, 1680) , 30-m
S4 Landsat-8 OLI Thick Cloud Removal Dallas, Texas, United States 2014-01 to 2014-11 (2040, 2040) , 30-m
  S5 Landsat-7 ETM+ Real ETM+ SLC-off Kern, California, United States 2011-09 to 2011-11 (1720, 2040) , 30-m
S6 Landsat-5 TM Real Thick Cloud Removal Urumqi, Xinjiang, China 2006-06 to 2009-08 (1440, 1440) , 30-m
IV Experiment
IV-A Experiment Settings
To assess the performance of the proposed algorithm, we conducted experiments on both simulated and real-world datasets. Additionally, we investigated the effectiveness of each core component of the proposed method.
IV-A1 Datasets
Different datasets were tailored for different missing data reconstruction tasks. For instance, for simulated ETM+ SLC-off and thick cloud removal tasks, we selected two research areas each for model training and testing. For real-world SLC-off and thick cloud removal, we chose one research area each and evaluated the trained model within them. All research areas comprised multiple images captured at different times. We employed a sliding window approach with a stride of 60 and a size of 120$\times$120 to augment the sample amount, and then split them into training and testing sets. Specifically, Fig. 6 illustrates the spatial distribution and examples of each research area, while Table I provides detailed information on their data sources, tasks, time ranges, etc.
IV-A2 Comparison Algorithms and Evaluation Metrics
For the ETM+ SLC-off and thick cloud cover problems, we compared our method with LLHM [21], WLR [24], and STS-CNN [37]. These algorithms take a target image and a single temporal auxiliary image as input and synthesise a single reconstructed image, thus lacking the ability to handle long temporal auxiliary images, i.e., Dual-In Single-Out. In contrast, our proposed method leverages multi-temporal images with missing data as input and outputs reconstructed multi-temporal images in parallel, i.e., Multi-In Multi-Out. The simulated experiments utilized the mean peak signal-to-noise ratio (mPSNR) and the mean structural similarity index (mSSIM) for all spectral bands [6] as evaluation metrics. Furthermore, we repeated each experiment 5 times and used the average results as the final evaluation metric.
IV-A3 Implement Detail
@@
For deep learning-based models, we utilized the Adam optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) for training, and set the batch size to 8. The initial learning rate was set to $4\times 10^{-4}$, with a decay schedule of every 100 epochs, and an early stopping strategy of 30 epochs on the validation set to prevent overfitting. The proposed MS2TAN was implemented with PyTorch 1.12 framework and trained with a NVIDIA GeForce RTX3090 24GB GPU on a Ubuntu 20.04 environment. Additionally, the test code of MS2TAN is openly available on the Github at https://github.com/zzaiyan/MS2TAN.
IV-B Simulated Experiment
Figure 7: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S1. (a) Ground truth (March 19, 2002). (b) Simulated ETM+ SLC-off. (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
Figure 8: Experimental results of Simulated ETM+ SLC-off Restoration based on Research Area S2. (a) Ground truth (April 15, 2003). (b) Simulated ETM+ SLC-off. (c) LLHM+LPRM. (d) WLR+LPRM. (e) STS-CNN. (f) Ours.
IV-B1 Simulated ETM+ SLC-off Restoration Experiment

### v2

@@ III-B3 Positional Encoding
In the Transformer, positional encoding is added to the input sequences to introduce positional semantics, as depicted in Eq. (12). Here, $pos=t\times N+n\in\left[0,~{}TN-1\right]$ represents the spatiotemporal joint coordinate derived from the temporal index $t\in\left[0,~{}T-1\right]$ and spatial index $n\in\left[0,~{}N-1\right]$, where $N$ denotes the total number of patches in a single image. The function $\operatorname{PosEnc}(pos,dim)$ indicates the value of the $dim$-th dimension of the positional encoding for the $pos$-th patch in the image sequence.
@@
After positional encoding, the input sequence will undergo $L$ layers of MSTA to explore deeper spatiotemporal features. Finally, the overall expression of the MFE module is shown in Eq. (13), where the symbol L denotes stacking $L$ layers, and $\mathrm{MFE}()$ represents the MSTA-based Feature Extractor.
@@ III-C Multi-scale Restoration Network
The MS2TAN (Multi-Scale Masked Spatial-Temporal Attention Network) is a multi-scale restoration network composed of $S$ residual-connected restoration modules and observed value replacement, as illustrated in Fig. 3. Each module consists of Patch Embedding, MFE, and Patch Unembedding. These restoration modules can capture the spatiotemporal correlations of pixels at different scales to predict missing values and ultimately replace observed values as outputs.
@@ III-C1 Embedding and Unembedding
Following the approach of ViT, each image is partitioned into a series of patches, with the size of each patch determined by the hyperparameter $P$. For each temporal image, $N=HW/P^{2}$ patches are obtained. These patches are then aggregated and reorganized to form a patch sequence of length $TN$, with each patch represented as a $CP^{2}$-dimensional vector. Equation (14) delineates the processing of the $t$-th temporal image $X_{\mathrm{in}}^{t}$.
$\begin{gathered}A^{t}=\left[p_{1}^{t},p_{2}^{t},\cdots,p_{N}^{t}\right]=\operatorname{Patchify}(x_{\mathrm{in}}^{t},P)\\ \text{where}\ x_{\mathrm{in}}^{t}\in\mathbb{R}^{C\times H\times W}\ \text{and}\ A^{t}\in\mathbb{R}^{N\times CP^{2}}\end{gathered}$ (14)
Next, a linear layer is utilized to project the vectors corresponding to each patch into a high-dimensional representation space. The images from different time steps are then aggregated to obtain the complete token sequence $E$ :
@@
The process of transforming $X_{\mathrm{in}}$ and $M$ step by step to obtain $E$ as described above is referred to as Embedding, which can be represented by Eq. (16). Conversely, if this operation is reversed, i.e., $E$ is restored to a patch sequence and then rearranged into an image, it is referred to as Unembedding, as shown in Eq. (17). Fig. 4 illustrates the processes of Embedding and Unembedding on images.
$E=\mathrm{Embedding}\left(X_{\mathrm{in}},P\right)$ (16)
@@ III-C2 MS2TAN
MS2TAN consists of components such as Embedding, MFE, and Unembedding, where the patch size $P$ is an important hyperparameter, and the embedding scale of the $i$-th layer is denoted as $P^{\left(i\right)}$. The process of obtaining the embedding vector $\alpha$ from the input is described in Eq. (18). Subsequently, MFE is utilized to mine deep spatiotemporal correlation features $\beta$, as shown in Eq. (19). Finally, the features are decoded, unembedded back into image form, and added to the input, resulting in an intermediate result as depicted in Eq. (20), where $\widetilde{Y}^{(i)}$ represents the $i$-th intermediate result.
$\alpha=\mathrm{Embedding}\left(\widetilde{Y}^{(i-1)},P^{\left(i\right)}\right)$ (18)
@@
Let $\widetilde{Y}^{(0)}=X_{\mathrm{in}}$, and sequentially obtain $S$ intermediate results of the reconstruction. Eventually, the $S$-th result is taken as the output of MS2TAN, as shown in Eq. (21).
@@ III-C3 Observed Value Replacement
For the original reconstruction results of this network, denoted as $\widetilde{Y}$, the observed values are replaced with their non-missing parts to obtain the final result $\widetilde{Y}_{\mathrm{out}}$, as shown in Eq. (3).
III-C4 Multi-scale Restoration Strategy
Common ViT-based models use a fixed patch size, which has been shown to be inefficient in dense prediction tasks like image segmentation and restoration [46]. In such tasks, smaller patch sizes often achieve better performance, but also come with greater computational costs. In order to balance the effect and efficiency, we apply a multi-scale restoration strategy, applying a larger patch size at a scale close to the input to obtain a coarse result, and applying a smaller patch size at a scale close to the output to obtain a fine result.
@@ III-D1 Pixel-wise Loss
Pixel-wise loss disregards the overall integrity of the image, treating the image as a collection of pixels, and comparing pixel by pixel to generate the image against the target image. It serves as the foundational loss for image reconstruction tasks. Equation (22) illustrates its calculation process. The pixel-wise loss calculated here includes both missing and observed parts. The reduction of the loss in the observed part does not directly improve the performance of the model. However, it does smooth the model output, making the model training process more stable, and hence is also included.
@@ III-D2 Structural Loss
Pixel-wise loss is commonly employed in various time-series and visual tasks but fails to consider the correlation between pixels. Structural loss uses structural similarity [47] to measure the difference from the target image to optimise the visual consistency of the reconstruction results in terms of structure, contrast and luminance, and has been shown to have better performance in image reconstruction tasks [48]. Equation (23) specifies the computation process of structural loss, where $\mu_{\eta}$ and $\mu_{y}$ denote the means of $\eta$ and $y$ respectively, $\sigma_{\eta}$ and $\sigma_{y}$ denote the variances of $\eta$ and $y$ respectively, $\sigma_{\eta y}$ denotes the covariance between $\eta$ and $y$, and $C_{1}$ and $C_{2}$ are constants.
@@ III-D3 Perceptual Loss
Perceptual loss [12] is widely used in tasks like super-resolution and image generation, where pre-trained CNNs are used as the feature network to extract perceptual features such as texture details, image style, etc. Specifically, we use a VGG-16 [49] with the fully connected layer removed as the feature network and the L2 loss of the feature map differences as the final loss. The calculation of perceptual loss is shown in Eq. (24), where $\psi\left(\right)$ denotes the feature network and $d_{\text{f}}$ denotes the feature dimension.
${\mathcal{L}}_{\mathrm{perceptual}}=\frac{1}{d_{\text{f}}}\left|\left|\psi\left(\eta\right)-\psi\left(y\right)\right|\right|_{2}^{2}$ (24)
@@ III-D4 Multi-Objective Loss Function
For each scale of MS2TAN, the expression of the loss function $\mathcal{L}^{\left(i\right)}$ corresponding to its intermediate output $\widetilde{Y}^{(i)}$ is shown in Eq. (25), where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ denote the weights of the pixel-wise loss, structural loss and perceptual loss respectively. The multi-objective loss function $\mathcal{L}$ considers the intermediate results of each scale, using their mean as the overall loss, as shown in Eq. (26).
@@
$\mathcal{L}^{\left(i\right)}\left(\eta,y\right)={\lambda_{1}}\mathcal{L}_{\text{pixel-wise}}+{\lambda_{2}}\mathcal{L}_{\mathrm{struct}}+{\lambda_{3}}\mathcal{L}_{\mathrm{perceptual}}$ (25)
$\mathcal{L}=\frac{1}{S}\sum_{i=1}^{S}\mathcal{L}^{\left(i\right)}\left(\widetilde{Y}^{(i)},Y\right)$ (26)
IV Experiments
IV-A Settings
To assess the performance of the proposed model in different scenarios, we conducted quantitative and qualitative experiments on two datasets and compared it with some mainstream methods. In addition, we conduct ablation experiments on the key innovations and discuss the balance between effectiveness and efficiency for models with different sizes.
IV-A1 Compared Algorithms
Based on the type of inputs and outputs, common image restoration methods can be categorized into three main types: single-input-single-output (SISO), dual-input-single-output (DISO), and multi-input-multi-output (MIMO). The SISO algorithms are the simplest type, but they struggle to achieve excellent performance, so we do not compare this category with ours. The DISO algorithm takes the target image and a temporal reference image as input and synthesizes a single reconstructed image. For these methods, we use the less missing temporal neighbor of the target image as the reference. The DISO methods we compare include simple replacement (replace missing values by copying data from reference image), LLHM [21], WLR [24], and STS-CNN [37]. The MIMO method utilizes a multi-temporal image containing missing data as input and outputs a reconstructed multi-temporal image, allowing the model to fully take into account temporal semantics. Compared methods of this type include last padding (Last), nearest neighbor padding (Nearest), and linear interpolation (Linear), which are commonly used in cartography [50], as well as the U-TILISE [40].
IV-A2 Evaluation Metrics
Through the sliding windows approach, we reconstructed and assessed data gaps across all time steps to facilitate a more comprehensive analysis. The quantitative experiments utilized mean absolute error (MAE), spectral angle mapper (SAM) [51], mean peak signal-to-noise ratio (PSNR), and mean structural similarity (SSIM) [47] as evaluation metrics. In addition, for real data gaps where ground truth is missing, we clearly show the comparison of restoration results for qualitative assessment.
IV-A3 Implement Details
@@
For MS2TAN, we utilized the Adam optimizer with parameters ($\beta_{1}=0.9$, $\beta_{2}=0.999$) for training, and set the batch size to 8. The initial learning rate was set to $4\times 10^{-4}$, with a decay schedule of every 100 epochs, and an early stopping strategy of 30 epochs on the validation set to prevent overfitting. For the other models we compared, we conducted our experiments while following the original authors’ training settings and environment. The proposed MS2TAN was implemented with PyTorch 1.12 framework and trained with a NVIDIA GeForce RTX 3090 24GB GPU on a Ubuntu 20.04 environment. Additionally, the test code of MS2TAN is openly available on the Github at https://github.com/CUG-BEODL/MS2TAN.
IV-B Datasets
We conducted experiments on two datasets, Landsat and EarthNet2021 [52]. Their details are described in Table VI.
TABLE I: Quantitative Evaluation Results under Dual-Temporal Inputs Compared to the DISO Methods, using the Landsat and EarthNet2021 datasets. The results for the best-performing method for each metric are highlighted in bold.
  Method Type Landsat EarthNet2021
$\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
  Replace DISO 0.0210 2.81 32.67 0.8913 0.0146 3.10 34.04 0.9452
LLHM [21] 0.0161 1.95 34.64 0.9104 0.0121 2.59 35.54 0.9566
WLR [24] 0.0110 1.33 36.88 0.9312 0.0113 2.41 36.08 0.9605
STS-CNN [37] 0.0091 1.12 37.79 0.9423 0.0102 2.18 37.07 0.9647
  MS2TAN (Ours) DIDO 0.0081 1.02 38.54 0.9503 0.0095 2.04 37.60 0.9671
Figure 5: Visual comparison results between MS2TAN and the DISO algorithms on the Landsat dataset, with dead pixels shown in black. Row 2 details the highlighted area in row 1. From left to right: (a) Input 1 (target). (b) Input 2 (reference). (c)-(g) restoration results of each method. (h) Ground truth.
IV-B1 Landsat
The Landsat dataset contains a large number of Top-of-Atmosphere (TOA) time-series captured by Landsat-5/7/8 satellites, with approximately uniform time intervals. The original images are cropped into small patches of size $120\times 120$ pixels with a spatial resolution of 30 meters. Each image includes 6 bands (Blue, Green, Red, NIR, SWIR1, and SWIR2), as well as the SLC-off or cloud mask provided in the QA band. In terms of spatial extent, we focus on sequences from multiple study regions in China and the United States.
IV-B2 EarthNet2021
The EarthNet2021 [52] dataset comprises over 28,000 Sentinel-2 TOA time-series observations from the European continent, with uniformly spaced time intervals. Each image consists of 4 bands (Blue, Green, Red and NIR), at a size of $128\times 128$ pixels, downsampled to a spatial resolution of 20-meters, accompanied by cloud probability maps. We utilized 80% of sequences from the training split for training purposes, with the remaining set aside for validation. For testing, we used sequences from the iid test split.
IV-B3 Preprocessing
We follow the preprocessing protocol of EarthNet2021 to crop the values of the optical images to the range $\left[0,10~{}000\right]$ and then normalize them to the unit range $\left[0,1\right]$. To match the multi-scale patch division strategy of MS2TAN, we resize the input images to a common multiple of the patch size of the MSTA at different scales. In our experiments, we use three scales (patch size is 8, 10, and 12 respectively) which have a least common multiple of 120. We therefore centrally cropped the images in the EarthNet2021 dataset to $120\times 120$ pixels, in line with the Landsat dataset.
IV-C Results
TABLE II: Quantitative Evaluation Results under Multit-Temporal Inputs Compared to the MIMO Methods, using the Landsat and EarthNet2021 datasets. The results for the best-performing method for each metric are highlighted in bold.
  Method Type Landsat EarthNet2021
$\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM $\downarrow$ MAE $\downarrow$ SAM $\uparrow$ PSNR $\uparrow$ SSIM
  Last MIMO 0.0223 3.14 32.11 0.8862 0.0148 3.17 33.68 0.9439
Nearest 0.0198 2.54 33.28 0.8957 0.0128 2.74 35.04 0.9534
Linear 0.0103 1.26 37.20 0.9348 0.0110 2.35 36.48 0.9620
U-TILISE [40] 0.0082 1.05 38.45 0.9491 0.0086 1.87 38.29 0.9702
MS2TAN (Ours) 0.0074 0.96 39.01 0.9552 0.0078 1.71 38.91 0.9728
Figure 6: Visual comparison results between MS2TAN and the MIMO algorithms on the EarthNet2021 dataset, with thick cloud shown in white. Each row shows a time-ordered image sequence. From top to bottom: input sequence (row 1), restoration results of each method (rows 2-6), and ground truth (row 7).
To quantitatively compare restoration results, we evaluated our proposed approach as well as mainstream methods on synthetic data gaps. Specifically, we performed comparison experiments under two different conditions: dual-temporal inputs (vs. the DISO algorithms), and multi-temporal inputs (vs. the MIMO algorithms). The quantitative evaluation results on the Landsat and EarthNet2021 datasets are shown in Table I and II, and visual eaxamples are illustrated in Figs. 5 and 6. Furthermore, Figs. 7 and 8 shows the restoration results in real data gaps to validate the generalisation performance of MS2TAN.
IV-C1 Comparison under Dual-Temporal Inputs